OpenStudy (anonymous):

When heated, KClO3 decomposes into KCl and O2. KClO3--> 2KCl+ 3O If this reaction produced 62.6 g of KCl, how much O2 was produced (in grams)?

OpenStudy (aaronq):

theres something wrong with your equation: O wouldn't exist alone, it would exist as O2

OpenStudy (anonymous):

2KClO3 --> 2KCl + 3O2 and i know that the KCl is 60.2 g but I don't know how to figure the amount of O in grams

OpenStudy (aaronq):

okay so you balanced it now find out how many moles of KCl you have, then use that to relate it to the number of moles of O2 based on the coefficients

OpenStudy (anonymous):

would it be3.23 mol? or would i multiply that by two to get 6.46 since there are two KCl

OpenStudy (aaronq):

set up a ratio \[x mol / 3 O2 = 0.8 mol/ 2 KCl \]

OpenStudy (anonymous):

511.3 g of O2?

OpenStudy (anonymous):

that number seems a little too high

OpenStudy (aaronq):

x= (0.8(3))/2= 1.2 moles m= n x MW = (1.2mol)(32 g/mol) =38.4 g

OpenStudy (aaronq):

sorry that wasn't 302 it was 3 Oxygen2

OpenStudy (anonymous):

arent they the same thing?

OpenStudy (aaronq):

i think you mistook \[3 O _{2} \] by the number three hundred and two

OpenStudy (anonymous):

ohhhh okay. so to determine how much O\[O _{2}\]

OpenStudy (anonymous):

what should i do? if 56.7 g of KCl is produced?

OpenStudy (anonymous):

i dont understand how you got the 0.8 earlier

OpenStudy (aaronq):

number of moles of KCl = mass KCl / molar mass of KCl n=m/MW= (62.6 g) / (74.55 g/mol) = 0.8397 moles (sorry it's 0.84 moles not 0.8 , i didn't have a calculator so i just did it in my head)

OpenStudy (aaronq):

if you had 56.7 g use n=m/MW to get the number of moles

OpenStudy (anonymous):

ok so the number of moles is .76 for KCl now i get the moles for \[O _{2}\]

OpenStudy (anonymous):

which would be 31.98, right?

OpenStudy (anonymous):

my final answer is 36.48 g of \[O _{2}\]

OpenStudy (aaronq):

thats right dude

OpenStudy (anonymous):

thanks so much!!!

OpenStudy (aaronq):

no prob !

OpenStudy (anonymous):

would you mind helping with one more problem? :)